# Uncertainty Quantification and Confidence Calibration in Large Language Models: A Survey

### v1

@@ 3.4.2. Multiple rounds generation
Token-Level Entropy. Token-level entropy quantifies uncertainty in LLMs by analyzing the probability distribution of generated tokens across multiple samples. A confident model assigns high probability to a specific token, resulting in low entropy, while uncertain predictions distribute probability across multiple tokens, leading to higher entropy.
@@
Multiple responses are generated for the same input to estimate token-level entropy, and the entropy of the token probability distribution is computed. For example, predictive entropy (Kadavath et al., 2022) can also be applied to multiple response settings and shows a better uncertainty quality based on multiple outputs variability. Similarly, SAR (Duan et al., 2024) could also be applied to multiple responses. Another study refines this method with Monte Carlo-based approximations (Malinin and Gales, 2021). It focuses on how probability distributions evolve across tokens during autoregressive generation. There are two main approaches: one that directly estimates uncertainty by averaging entropy across multiple sampled outputs and another that decomposes sequence-level uncertainty into token-level contributions using a structured entropy approximation.
@@
Conformal Prediction. Conformal Prediction (CP) is a statistical framework that provides formal coverage guarantees for uncertainty estimation in LLMs. It has been widely adopted in recent research due to its distribution-free properties, making it suitable for both black-box and white-box models.
@@
In the black-box setting, where model internals are inaccessible, CP estimates uncertainty using response frequency, semantic similarity, or self-consistency. One study proposes a method tailored for API-only LLMs (Su et al., 2024), using frequency-based sampling combined with normalized entropy (NE) and semantic similarity (SS) to define nonconformity scores. Another black-box CP method introduces a self-consistency-based uncertainty measure (Wang et al., 2024a), which clusters sampled generations and selects a representative response to construct prediction sets with correctness guarantees, making it particularly effective for open-ended NLG tasks.
@@
On the other hand, white-box CP methods use logits, internal activations, and calibration techniques for more refined uncertainty estimation. One study proposes Conformal Language Modeling (Quach et al., 2023), which integrates CP with autoregressive text generation by dynamically calibrating a stopping rule to ensure at least one response in the generated set is statistically valid. Another work adapts CP for multiple-choice QA (Kumar et al., 2023), using model confidence scores to calibrate prediction sets, ensuring coverage with minimal set size. A more advanced technique, conditional CP (Cherian et al., 2025), dynamically adjusts coverage guarantees based on the difficulty of the input, optimizing prediction set size while maintaining reliability.
@@
Consistency-Based Methods. Consistency-based uncertainty estimation methods analyze the agreement between multiple generated responses from an LLM to determine uncertainty. The underlying assumption is that if the model is confident, its responses should be consistent, while high variability among responses suggests uncertainty. One approach leverages Jaccard similarity (Lin et al., 2023), which measures the overlap between words in different generations. This method evaluates the deviation from self-consistency, where a high Jaccard similarity across generations implies low uncertainty.
@@
However, in natural language generation (NLG) tasks, word-level similarity alone is insufficient, as different responses can convey the same meaning using different phrasing. To address this problem, some methods incorporate external models to assess semantic similarity rather than relying solely on lexical overlap.
@@ 3.4.3. Multiple Rounds Generation with External Models
Semantic-based uncertainty estimation methods extend multiple rounds of generation by incorporating external models to assess the consistency of generated responses beyond lexical similarity.
Semantic Entropy. Semantic Entropy (SE) (Kuhn et al., 2023), refines uncertainty estimation by clustering generated responses based on semantic equivalence. This approach uses a Natural Language Inference (NLI) model to determine entailment relationships among responses, grouping them into meaning-preserving clusters. Instead of calculating entropy over individual responses, SE computes entropy over these clusters. Another method, Kernel Language Entropy (KLE), applies a kernel-based framework to quantify semantic uncertainty (Nikitin et al., 2024). To enhance the performance of semantic entropy, KLE represents them in a semantic space using positive semidefinite kernels. By computing von Neumann entropy over these response distributions, KLE provides an even more fine-grained measure of uncertainty that considers nuanced semantic variations.
Semantic Similarity. Semantic similarity uncertainty methods use Natural Language Inference (NLI) models to measure the semantic relationships between multiple generated responses from an LLM (Lin et al., 2023). Instead of relying on lexical overlap, these approaches construct a similarity matrix based on entailment and contradiction scores between generated outputs. A confident model produces responses with high internal consistency, while greater semantic dispersion in the similarity matrix indicates higher uncertainty.
To quantify this dispersion, graph-based spectral metrics are applied. Eccentricity (Ecc) measures the spread of response variability, eigenvalue-based measures (Eig) analyze the spectral properties of the similarity matrix to detect uncertainty, and degree (Deg) evaluates response connectivity, with higher degrees indicating stronger confidence. Chen et al. (2025) proposes MD-UQ, which enhanced the performance of graph-based spectral metrics by using an additional knowledge dimension input extracted from an auxiliary LLM and tensor decomposition while Da et al. (2024a) also enhanced the performance by analyzing the directed graph instead of undirected graph originally from Lin et al. (2023).
@@ 4.1. Benchmark Datasets
Consistency & Ambiguity ParaRel (Elazar et al., 2021), AmbigQA (Min et al., 2020b), AmbigInst (Hou et al., 2024)
Table 3. Categorization of datasets and benchmarks.
Datasets used in previous studies can be organized into several categories based on their focus. Reading comprehension benchmarks include CoQA for conversational Q&A, RACE for general reading comprehension, TriviaQA for fact-based questions, CosmosQA for contextual understanding, SQuAD for question-answering on passages, and HotpotQA for multi-hop reasoning. Reasoning and math benchmarks include HotpotQA and StrategyQA, which test multi-hop reasoning, GSM8K for solving math problems, and CalibratedMath, designed to evaluate confidence expression in arithmetic. Factuality evaluation draws on datasets such as TruthfulQA for addressing common misconceptions, FEVER for claim verification, and HaluEval for detecting hallucinations, and annotated FActScore dataset for evaluating the factuality of long-form text generated by LLMs (Zhang et al., 2024a). Several general knowledge benchmarks have been adapted for uncertainty quantification, such as MMLU for a wide range of subjects, GPQA for multiple-choice questions in physical sciences, and HellaSwag for common-sense reasoning through sentence completion. These benchmarks can be adapted for UQ because the task reduces to a binary classification problem—determining whether the model is confident or uncertain. The structured nature of these benchmarks allows for clear evaluation of the model’s confidence in its predictions.
Additional categories include consistency benchmarks such as ParaRel, which tests semantic consistency across 328 paraphrases for 38 relations, and datasets like AmbigQA and AmbigInst, which feature inherent ambiguities. Ambiguity datasets are useful in uncertainty quantification (UQ) evaluation because they introduce aleatoric uncertainty by highlighting cases where multiple plausible interpretations exist, helping to assess how well models distinguish between data-driven randomness and model-based uncertainty. These datasets enable a more precise decomposition of uncertainty into aleatoric and epistemic components, improving model reliability and interpretability.
Recently, there have been efforts to introduce dedicated uncertainty quantification (UQ) benchmarks for large language models (LLMs). Yang et al. (2024) introduced MAQA, a dataset specifically designed to evaluate epistemic uncertainty in language models. Fadeeva et al. (2023) created LM-Polygraph, which was later adopted as a comprehensive uncertainty benchmark by Vashurin et al. (2024). Ye et al. (2025) developed a benchmark using conformal prediction across five common NLP tasks. These contributions represent specialized datasets explicitly designed to assess UQ capabilities in LLMs, rather than adapting existing general-purpose benchmarks. Overall, we show the categorization of datasets and benchmarks for UQ in Table 3.
@@ 4.2. Evaluation Metrics
Uncertainty quantification (UQ) is often evaluated as a binary classification task, the rationale being that high uncertainty should correspond to low expected accuracy. This is typically modeled by assigning a binary label to each response with a correctness function and using the uncertainty estimates to predict the label. AUROC (Area Under the Receiver Operating Characteristic curve), which measures how effectively the uncertainty score separates correct from incorrect responses, is often used. With values ranging from 0 to 1, higher AUROCs indicate better performance. Responses with confidence above the threshold are classified as predicted positives, while those below are treated as predicted negatives. Many prior studies use AUROC to evaluate how well uncertainty score discriminates correct from incorrect predictions (Chen and Mueller, 2024; Lin et al., 2023; Kuhn et al., 2023; Xiong et al., 2024a; Liu et al., 2024c).
Similarly, AUPRC (Area Under the Precision-Recall Curve) and AUARC (Area Under the Accuracy-Rejection Curve) (Nadeem et al., 2009) also offer further insights into uncertainty quantification. AUPRC measures how well the uncertainty score separates correct from incorrect responses (Ling et al., 2024), while AUARC assesses how effectively the uncertainty measure aids in selecting accurate responses by determining which uncertain questions to reject (Lin et al., 2024).
@@
In the context of NLG where the correctness label is hard to obtain, researchers also compute heuristic-based (fuzzy matching) metrics such as BLEU (Papineni et al., 2002) and ROUGE (Kuhn et al., 2023) between the generated text and the reference output(s) to gauge the quality. However, these metrics often fail to capture semantic fidelity or factual correctness. Consequently, many researchers are increasingly turning to LLM-as-a-judge evaluations, wherein a large language model (e.g., GPT-4) is prompted to assess text quality or correctness. This approach can capture nuanced aspects like coherence, style, and factuality, but also introduces risks of bias and inconsistency. Human annotation, however, is expensive and is often limited to a small scale (Zhang et al., 2024a; Kuhn et al., 2023).
@@
Apart from the binary classification framework, there are also multiple evaluation methods designed for specific treatment of uncertainty, sometimes qualitative. For example, focusing on decomposing aleatoric and epistemic uncertainty, Hou et al. (2024) evaluates only the aleatoric part by using AmbigQA (Min et al., 2020b), as high ambiguity questions should incur higher aleatoric uncertainty (whereas math questions, for examples, might have lower). The evaluation in Giulianelli et al. (2023), on the other hand, is a comparison between the variability of human production (generation) with that of the LM. With an emphasis on UQ for longer generations, Zhang et al. (2024a) compare the uncertainty estimate against FActScore (Min et al., 2023), as the “correctness” of a long paragraph could be ill-defined or ambiguous.
5. Applications in LLMs
@@
LLMs are increasingly applied in diverse domains, offering flexibility and reasoning capabilities (Da et al., 2025b). However, uncertainty quantification (UQ) is crucial for ensuring their reliability, particularly in high-stakes applications like Education (Lee et al., 2024) and healthcare (Savage et al., 2025; Da et al., 2024d). This section will introduce the applications that integrate the UQ of LLMs from different domains.
@@
Robotics. LLM-based robotic planning suffers from ambiguity and hallucinations, motivating the need for uncertainty quantification in the planning loop. For example, closed-loop planners (Zheng et al., 2024) employs an uncertainty-based failure detector to continuously assess and adjust plans in real-time, while non-parametric UQ method (Tsai et al., 2024) use an efficient querying strategy to improve reliability. In addition, ’LAP’ (Mullen Jr and Manocha, 2024) integrates action feasibility checks to align the LLM’s confidence with real-world constraints, improving success rates from approximately 70% to 80%. Similarly, adaptive skill selection (Ong et al., 2024) dynamically adjusts thresholds for alternative paths, achieving higher success rates in the ALFRED domain. Finally, introspective planning (Liang et al., 2024b) enables LLMs to self-assess uncertainty, enhancing safety and human-robot collaboration.
@@
Transportation. Preliminary research explores how LLMs can enhance transportation systems (Da et al., 2024c). For example, LLM inference has been used to bridge the sim-to-real gap in Traffic Signal Control policies (Da et al., 2023b, a, 2024b), and LLM agents have been shown to help smooth mixed-autonomy traffic (Yao et al., 2024). However, both cases reveal the potential risk posed by hallucination. A few works have investigated the uncertainty measure while using the LLMs as in (de Zarzà et al., 2023); it tries to link the use of visual language models (VLMs) with deep probabilistic programming for uncertainty quantification while conducting multimodal traffic accident forecasting tasks and (Da et al., 2024a) shows a general solution by modeling the responses as a directed-graph and captures the response uncertainty. Further, UQ exploration anchored on traffic-related LLM use is expected.
@@
Healthcare. In healthcare, LLMs (VLMs) can be good reference for diagnosis, but uncertainty is a critical dimension that should be output together with the generation for more reliable treatment plans. Such as (Savage et al., 2024) evaluates the ability of uncertainty metrics to quantify LLM confidence when performing diagnosis and treatment selection tasks by assessing the properties of discrimination and calibration.  (Chen et al., 2024) first quantify uncertainty in white-box models with access to model parameters and output logits, and by this, it reveals that an effective reduction of model uncertainty can be achieved by using the proposed multi-tasking and ensemble methods in EHRs. With these findings, however, as  (Wu et al., 2024) benchmarks popular uncertain quantification methods with different model sizes on medical question-answering datasets, authors find that the challenge of UQ for medical applications is still severe.
Education. Many LLM-related products are popular to guide the thinking of students in AI for Education research (Zhang et al., 2024e). However, it is critical to ensure the trustworthiness of LLMs if to participate in a person’s thinking logic learning process (Razafinirina et al., 2024). The work in (Tziakouri, 2024) discusses the solutions of black-box models UQ evaluation, including Graph Laplacian Eigenvalue Sum, Lexical Similarity, and their reliability. (Da et al., 2025a) proposes a method that elicits the reasoning topology of the LLM answering process, which can guide the step-by-step analysis of the LLM’s reasoning steps before applying it to the student reference guide.
Conclusion. There are many different fields that discuss the need for uncertainty quantification during the LLMs execution process, and the above-discussed ones are only major aspects for now, the multi-agent energy management, operation research, etc, all employ LLMs and would require such consideration respectively.

### v2

@@ 3.4.2. Multiple rounds generation
$\bullet$ Token-Level Entropy. Token-level entropy quantifies uncertainty in LLMs by analyzing the probability distribution of generated tokens across multiple samples. A confident model assigns high probability to a specific token, resulting in low entropy, while uncertain predictions distribute probability across multiple tokens, leading to higher entropy.
@@
Multiple responses are generated for the same input to estimate token-level entropy, and the entropy of the token probability distribution is computed. For example, predictive entropy (Kadavath et al., 2022) can also be applied to multiple response settings and shows a better uncertainty quality based on the variability of multiple outputs. Similarly, SAR (Duan et al., 2024) could also be applied to multiple responses. (Malinin and Gales, 2021) extends with Monte Carlo-based approximations and focuses on how probability distributions evolve across tokens during autoregressive generation. There are two main approaches to get the final uncertainty: one averages entropy across multiple sampled outputs, and the other decomposes sequence-level uncertainty into token-level contributions using entropy approximation.
@@
$\bullet$ Conformal Prediction. Conformal Prediction (CP) (Shafer and Vovk, 2008) is a statistical framework that provides formal coverage guarantees for uncertainty estimation in LLMs. Its distribution-free properties make it suitable for both black-box and white-box models.
@@
In the black-box setting, where model internals are inaccessible, CP estimates uncertainty using response frequency, semantic similarity, or self-consistency. One study proposes a method tailored for API-only LLMs (Su et al., 2024), using frequency-based sampling combined with normalized entropy and semantic similarity to define nonconformity scores. Another black-box CP method introduces a self-consistency-based uncertainty measure (Wang et al., 2024a), which clusters sampled generations and selects a representative response to construct prediction sets with correctness guarantees, making it particularly effective for open-ended NLG tasks.
@@
On the other hand, white-box CP methods use logits, internal activations, and calibration techniques for more refined uncertainty estimation. One study proposes Conformal Language Modeling (Quach et al., [n. d.]), which integrates CP with autoregressive text generation by dynamically calibrating a stopping rule to ensure at least one response in the generated set is statistically valid. Another work adapts CP for multiple-choice QA (Kumar et al., 2023), using model confidence scores to calibrate prediction sets, ensuring coverage with minimal set size. A more advanced technique, conditional CP (Cherian et al., 2025), dynamically adjusts coverage guarantees based on the difficulty of the input, optimizing prediction set size while maintaining reliability.
@@
$\bullet$ Consistency-Based Methods. Consistency-based uncertainty estimation methods analyze the agreement between multiple generated responses from an LLM to determine uncertainty. The underlying assumption is that if the model is confident, its responses should be consistent, while high variability among responses suggests uncertainty.  (Lin et al., 2023) measures the overlap between words through Jaccard similarity in different generations. This method evaluates the deviation from self-consistency, where a high Jaccard similarity across generations implies low uncertainty.
@@
However, word-level similarity alone is insufficient, as different responses can convey the same meaning using different phrasing. Moreover, the generated response might include long reasoning steps that require detailed analysis (Golovneva et al., [n. d.]). To address this problem, some methods incorporate external models to assess semantic similarity rather than relying solely on lexical overlap.
@@ 3.4.3. Multiple Rounds Generation with External Models
Semantic-based uncertainty estimation methods expand multiple generation approaches by incorporating external models, such as Natural Language Inference (NLI) or pretrained language models, to evaluate the semantic relationships among generated responses beyond surface-level similarity.
$\bullet$ Distribution-based entropy methods quantify uncertainty by modeling the distribution of generated responses in a semantic space. Semantic Entropy (SE) (Kuhn et al., 2023) refines uncertainty estimation by clustering generated responses based on semantic equivalence. This approach uses an NLI model to determine entailment relationships among responses, grouping them into meaning-preserving clusters. Instead of calculating entropy over individual responses, SE computes entropy over these clusters. Kernel Language Entropy (KLE) (Nikitin et al., 2024) takes a different approach by avoiding explicit clustering. Instead, it embeds the responses in a semantic space using a positive semidefinite kernel function. By computing von Neumann entropy over these response distributions, KLE provides an even more fine-grained measure of uncertainty that considers nuanced semantic variations.
$\bullet$ Pairwise similarity methods construct a pairwise semantic similarity matrix between responses and analyze its structural properties to estimate uncertainty. Methods like (Lin et al., 2023) use NLI models to score entailment and contradiction between every pair of generated outputs, forming a weighted similarity graph. A confident model yields semantically coherent responses with strong mutual agreement (high similarity), while inconsistent or ambiguous outputs lead to greater dispersion in the matrix. To quantify this dispersion, spectral graph metrics are applied: Eccentricity (Ecc) captures variability spread, Eigenvalue-based (Eig) measures assess global structure, and Degree (Deg) evaluates local consistency. Recent works further extend this by modeling the response similarity graph as directed (Da et al., 2024a) or multi-dimensional (Chen et al., 2025b), allowing for richer representation of semantic asymmetry or latent factors in uncertainty.
@@ 4.1. Benchmark Datasets
Datasets used in previous studies can be organized into several categories based on their focus. An overall summary of the categorization of datasets and benchmarks for UQ is shown in Table 3.
$\bullet$ Reading comprehension benchmarks include CoQA (Reddy et al., 2019) for conversational question answering tasks, RACE (Lai et al., 2017) for general reading comprehension, TriviaQA (Joshi et al., 2017) for fact-based questions, CosmosQA (Huang et al., 2019) for contextual understanding, SQuAD (Rajpurkar et al., 2016) for question answering on passages, and HotpotQA (Yang et al., 2018) for multi-hop reasoning.
$\bullet$ Reasoning and math benchmarks include HotpotQA (Yang et al., 2018) and StrategyQA (Geva et al., 2021), which test multi-hop reasoning, GSM8K (Cobbe et al., 2021) for solving math problems, and CalibratedMath (Lin et al., 2022a), designed to evaluate confidence expression in arithmetic. These benchmarks are helpful to evaluate the reasoning uncertainty.
$\bullet$ Factuality evaluation draws on datasets such as TruthfulQA (Lin et al., 2022b) for addressing common misconceptions, FEVER (Thorne et al., 2018) for claim verification, HaluEval (Li et al., 2023a) for detecting hallucinations, and an annotated FActScore (Min et al., 2023) dataset for evaluating the factuality of long-form text generated by LLMs.
$\bullet$ General knowledge benchmarks can be adapted for UQ to test the models’ general knowledge, such as MMLU (Hendrycks et al., 2020) for a wide range of subjects, GPQA (Rein et al., 2024) for multiple-choice questions in physical sciences, and HellaSwag (Zellers et al., 2019) for common-sense reasoning through sentence completion. These benchmarks can be adapted for UQ because the tasks can be reduced to a classification problem, determining whether the model is confident or uncertain. The structured nature of these benchmarks allows for clear evaluation of the model’s confidence in its predictions.
$\bullet$ Consistency and ambiguity are two additional kind of benchmarks for UQ. Consistency benchmarks such as ParaRel (Elazar et al., 2021) tests semantic consistency across 328 paraphrases for 38 relations, and datasets like AmbigQA and AmbigInst, which feature inherent ambiguities (Min et al., 2020; Hou et al., 2024; Chen et al., 2025a). Ambiguity datasets are useful in UQ evaluation because they introduce aleatoric uncertainty by highlighting cases where multiple plausible interpretations exist, helping to assess how well models distinguish between data-driven randomness and model-based uncertainty. These datasets enable a more precise decomposition of uncertainty into aleatoric and epistemic components, improving model reliability and interpretability.
@@
Consistency & Ambiguity ParaRel (Elazar et al., 2021), AmbigQA (Min et al., 2020), AmbigInst (Hou et al., 2024), Abg-SciQA (Chen et al., 2025a)
Table 3. Categorization of benchmarking datasets for UQ.
Recently, there have been efforts to develop UQ benchmarks for dedicated sources of uncertainty or specific methods. For example, MAQA (Yang et al., 2025) is a dataset specifically designed to evaluate epistemic uncertainty in language models; LM-Polygraph (Fadeeva et al., 2023) was later adopted as a comprehensive uncertainty benchmark (Vashurin et al., 2025). (Ye et al., 2025) developed a benchmark for conformal prediction methods. These contributions represent specialized datasets explicitly designed to assess UQ capabilities in LLMs, rather than adapting existing general-purpose benchmarks.
@@ 4.2. Evaluation Metrics
UQ is often evaluated from binary classification tasks, with the rationale being that high uncertainty should correspond to low expected accuracy. This is typically modeled by assigning a binary label to each response with a correctness function and using the uncertainty estimates to predict the label. AUROC (Area Under the Receiver Operating Characteristic curve), which measures how effectively the uncertainty score separates correct from incorrect responses, is often used. With values ranging from 0 to 1, higher AUROCs indicate better performance. Responses with confidence above the threshold are classified as predicted positives, while those below are treated as predicted negatives. Many prior studies use AUROC to evaluate how well the uncertainty score discriminates correct from incorrect predictions (Chen and Mueller, 2024; Kuhn et al., 2023; Xiong et al., 2024a; Liu et al., 2024c, 2025). Similarly, AUPRC (Area Under the Precision-Recall Curve) and AUARC (Area Under the Accuracy-Rejection Curve) (Nadeem et al., 2009) also offer further insights into UQ. AUPRC measures how well the uncertainty score separates correct from incorrect responses (Ling et al., 2024), while AUARC assesses how effectively the uncertainty measure aids in selecting accurate responses by determining which uncertain questions to reject (Lin et al., 2024).
@@
In the context of NLG where the correctness label is hard to obtain, researchers also compute heuristic-based fuzzy matching metrics such as BLEU (Papineni et al., 2002) and ROUGE (Kuhn et al., 2023) between the generated text and the reference output(s) to gauge the quality. However, these metrics often fail to capture semantic fidelity or factual correctness. Consequently, many researchers are increasingly turning to LLM-as-a-judge evaluations, wherein a large language model (e.g., GPT-4) is prompted to assess text quality or correctness. This approach can capture nuanced aspects like coherence, style, and factuality, but also introduces risks of bias and inconsistency. Human annotation, however, is expensive and is often limited to a small scale (Zhang et al., 2024b; Kuhn et al., 2023).
@@
Apart from the binary classification framework, there are also multiple evaluation methods designed for the specific treatment of uncertainty, sometimes qualitative. For example, focusing on decomposing aleatoric and epistemic uncertainty, (Hou et al., 2024) evaluates only the aleatoric part by using AmbigQA (Min et al., 2020), as high ambiguity questions should incur higher aleatoric uncertainty (whereas math questions, for examples, might have lower). The evaluation in (Giulianelli et al., 2023), on the other hand, is a comparison between the variability of human production (generation) with that of the LM. With an emphasis on UQ for longer generations, (Zhang et al., 2024b) compares the uncertainty estimate against FActScore (Min et al., 2023), as the “correctness” of a long paragraph could be ill-defined or ambiguous.
5. UQ Applications in LLMs
@@
LLMs are increasingly applied in diverse domains, offering flexibility and reasoning capabilities. However, UQ is crucial for ensuring their reliability, particularly in high-stakes applications. This section will introduce the applications that integrate the UQ of LLMs from some example domains. Many other fields like energy management, operations research, etc., employ LLMs and would require such discussions on the need for UQ as well.
@@
$\bullet$ Robotics. LLM-based robotic planning suffers from ambiguity and hallucinations, motivating the need for UQ in the planning loop. For example, closed-loop planners (Zheng et al., 2024) employ an uncertainty-based failure detector to continuously assess and adjust plans in real-time, while non-parametric UQ methods (Tsai et al., 2024) use an efficient querying strategy to improve reliability.  (Mullen Jr and Manocha, 2024) integrates action feasibility checks to align the LLM’s confidence with real-world constraints, improving success rates from approximately 70% to 80%. Similarly, (Ong et al., 2024) dynamically adjusts thresholds for alternative paths in adaptive skill selection, achieving higher success rates.  (Liang et al., 2024b) develops an introspective planning framework with LLMs self-assess their uncertainty to enhance safety and human-robot collaboration.
@@
$\bullet$ Transportation. Preliminary research explores how LLMs can enhance transportation systems (Da et al., 2024c; Yao et al., 2025; Da et al., 2023). For example, LLM inference has been used to bridge the sim-to-real gap in traffic signal control (Da et al., 2023, 2024b) and smooth mixed-autonomy traffic (Yao et al., 2025). However, both cases reveal the potential risk posed by hallucination. A few works have investigated the uncertainty measure while using the LLMs (de Zarzà et al., 2023), which tries to link the use of VLMs with deep probabilistic programming for UQ while conducting multimodal traffic accident forecasting tasks.
@@
$\bullet$ Healthcare. In healthcare, LLMs and VLMs can be good references for diagnosis, but uncertainty is a critical dimension that should be considered together with the generation of more reliable treatment plans (Savage et al., 2024). In (Chen et al., 2024), it quantifies uncertainty in a white-box setting, and reveals that an effective reduction of model uncertainty can be achieved by using the proposed multi-tasking and ensemble methods in EHRs. However, as (Wu et al., 2024) benchmarks popular uncertain quantification methods with different model sizes on medical question-answering datasets, the challenge of UQ for medical applications is still severe.